OpenStudy (anonymous):

If you toss 2 dice 10,000 times, how many times would you expect that " he sum of the faces equals 4" happens? Answer 733 550 333 833

OpenStudy (dan815):

833

OpenStudy (dan815):

explaination : 3/36 chance when you roll dice you will get a sub of 4 1-3 3-1 2-2 those are the 3 ways and all the possibilities of dice = 6*6 = 36 3/36 = 1/12 possibility outta 10000 rolls 10000/12 = 833.3333

OpenStudy (dan815):

sum not sub*

OpenStudy (anonymous):

http://www.weegy.com/?ConversationId=58523B29

OpenStudy (anonymous):

@ crazy girl weegys answer was not correct, (dan815) was correct thank you though,I appreciate you guys help.

OpenStudy (anonymous):

srry

OpenStudy (anonymous):

its okay I appreciate your help.

OpenStudy (dan815):

np :)